Calculate the rotational inertia of a meter stick, with mass 0.74 kg, about an axis perpendicular to the stick and located at the 34 cm mark. (Treat the stick as a thin rod.) kg·m2
Answer #1


monet of inertia I = Ic + m*r^2


c = momentof inertia about the cente rof mas = (1/12)*m*L^2


L = length of the stick = 1 m


m = mass of the stick = 0.74 kg

r = distance between center of mass and rotating axis = 50 - 34 = 16 cm =0.16 m


I = ((1/12)*0.74*1^2) + (0.74* 0.16^2)


I = 0.0806 kg m^2 <<<------------ANSWER